# Topologically Protected Surface Altermagnetism on Antiferromagnets

### v1

@@ C.1 Non-magnetic part
$H$ $=2t\sigma_{x}(\cos\tfrac{k_{x}+k_{y}}{2}+\cos\tfrac{k_{x}-k_{y}}{2})+2t_{\perp}\tau_{x}\cos\tfrac{k_{z}}{2}$ (27)
@@
$\bm{t}_{\tfrac{a}{2}\tfrac{c}{2}}$ $=I\{\tfrac{1}{2},0,\tfrac{1}{2}\}$ $H$ $=\sigma_{x}H\sigma_{x}$ (28)
$\bm{t}_{\tfrac{b}{2}}$ $=I\{0,\tfrac{1}{2},0\}$ $H$ $=\tau_{x}H\tau_{x}$ (29)
$\bm{t}_{\tfrac{a}{2}\tfrac{b}{2}\tfrac{c}{2}}$ $=I\{\tfrac{1}{2},\tfrac{1}{2},\tfrac{1}{2}\}$ $H$ $=\sigma_{x}\tau_{x}H\sigma_{x}\tau_{x}$ (30)
which can give rise to accidental constraints for the band structure, as we show below. They can be broken by further neighbor hoppings, see Fig. 5(b). If we add
$\delta H=$ $2(t_{1}+t_{2})\sigma_{x}\tau_{x}\left[\cos\tfrac{k_{x}+k_{y}}{2}+\cos\tfrac{k_{x}-k_{y}}{2}\right]\cos\tfrac{k_{z}}{2}+2(t_{1}-t_{2})\sigma_{y}\tau_{y}\left[\sin\tfrac{k_{x}+k_{y}}{2}+\sin\tfrac{k_{x}-k_{y}}{2}\right]\sin\tfrac{k_{z}}{2}$ (31)
@@
$\delta H$ $=2\sigma_{x}[(t_{3}+t_{4})\cos k_{z}(\cos\tfrac{k_{x}+k_{y}}{2}+\cos\tfrac{k_{x}-k_{y}}{2})]+2\sigma_{x}\tau_{z}[(t_{3}-t_{4})\sin k_{z}(\sin\tfrac{k_{x}+k_{y}}{2}-\sin\tfrac{k_{x}-k_{y}}{2})]$ (32)
@@
$C_{2b}^{\prime}=C_{2b}\{\tfrac{1}{2},\tfrac{1}{2},\tfrac{1}{2}\}*I\{0,\tfrac{1}{2},0\}$ $=C_{2b}\{\tfrac{1}{2},0,\tfrac{1}{2}\}=\sigma_{x}$ (33)
$C_{2b}^{\prime\prime}=C_{2b}\{\tfrac{1}{2},\tfrac{1}{2},\tfrac{1}{2}\}*I\{\tfrac{1}{2},\tfrac{1}{2},\tfrac{1}{2}\}$ $=C_{2b}\{0,0,0\}$ (34)
@@ C.2 Magnetic part
$H_{M}=$ $2\Delta_{0}\sigma_{y}s_{z}(\sin\tfrac{k_{x}+k_{y}}{2}-\sin\tfrac{k_{x}-k_{y}}{2})+2\Delta_{1}\sigma_{z}\tau_{y}s_{z}\sin\tfrac{k_{z}}{2}$ (35)
@@
$H_{M}^{\prime}=2\Delta_{2}\sigma_{y}\tau_{x}s_{z}(\sin\tfrac{k_{x}+k_{y}}{2}-\sin\tfrac{k_{x}-k_{y}}{2})\cos\tfrac{k_{z}}{2}]+2\Delta_{3}\sigma_{x}\tau_{y}s_{z}(\cos\tfrac{k_{x}+k_{y}}{2}-\cos\tfrac{k_{x}-k_{y}}{2})\sin\tfrac{k_{z}}{2}]$ (37)
@@
C.3 Symmetry breaking perturbations
@@
$\delta H$ $=\sigma_{z}$ $(-,+,-)$ (39)
$\delta H$ $=\tau_{z}$ $(-,-,+)$ (40)
$\delta H$ $=\sigma_{z}\tau_{z}$ $(+,-,-)$ (41)
@@
The perturbation $\tau_{z}$ therefore has the same symmetry as $u_{ab}$. However, it turns out that this perturbation is invariant under an accidental mirror $M_{b}^{\prime}=\sigma_{x}$ of the $k_{z}=0$ plane, and the nodal line is not gapped unless higher order, symmetry-allowed hoppings are included. Because of this, we consider another perturbation with the same symmetry $\sigma_{z}\tau_{z}\left[\cos\left(k_{x}+k_{y}\right)-\cos\left(k_{x}-k_{y}\right)\right]$, which is the hopping considered in the main text, and shown in Fig. 5 (d).
C.4 Spin-orbit coupling
Figure 6: Influence of SOC on the altermagnetic drumhead states. We consider each of the three SOC terms in Eq. 42, each time setting the other two to zero. (a-e) $\lambda_{1}=0.4$. Bulk 3D band structure (a), bulk spin-resolved spectral function (b,c), surface spin-resolved spectral function (d,e), and spin splitter angles $\alpha_{x},\alpha_{y}$ (f). In the presence of $\lambda_{1}$-term SOC the spin-space group notation breaks down such that $C_{2}||M_{a}]$ is not a symmetry anymore. (g-k) $\lambda_{2}=0.2$. Bulk 3D band structure (g), bulk spin-resolved spectral function (h,i), surface spin-resolved spectral function (j,k), and spin splitter angles $\alpha_{x},\alpha_{y}$ (l). The $\lambda_{2}$-term SOC gaps out the Dirac nodal line everywhere except at the BZ boundary, where 2 Dirac cones persist and spin remains a good quantum number. The topologically unprotected, altermagnetic surface state persists. (m-q) $\lambda_{3}=0.2$. Bulk 3D band structure (m), bulk spin-resolved spectral function (n,o), surface spin-resolved spectral function (p,q), and spin splitter angles $\alpha_{x},\alpha_{y}$ (r). The $\lambda_{3}$-term SOC gaps out the Dirac nodal line everywhere. The topologically unprotected, altermagnetic surface state persists.
Figure 7: Influence of SOC on the altermagnetic Fermi-arc state. We consider each of the three SOC terms in Eq. 42, each time setting the other two to zero. (a-e) $\lambda_{1}=0.4$. Bulk 3D band structure (a), bulk spin-resolved spectral function (b,c), surface spin-resolved spectral function (d,e), and spin splitter angles $\alpha_{x},\alpha_{y}$ (f). In the presence of $\lambda_{1}$-term SOC the spin-space group notation breaks down such that $C_{2}||M_{a}]$ is not a symmetry anymore. (g-k) $\lambda_{2}=0.2$. Bulk 3D band structure (g), bulk spin-resolved spectral function (h,i), surface spin-resolved spectral function (j,k), and spin splitter angles $\alpha_{x},\alpha_{y}$ (l). The $\lambda_{2}$-term SOC gaps out the Dirac cones along the $y$-axis.The topologically unprotected, altermagnetic surface state persists. (m-q) $\lambda_{3}=0.2$. Bulk 3D band structure (m), bulk spin-resolved spectral function (n,o), surface spin-resolved spectral function (p,q), and spin splitter angles $\alpha_{x},\alpha_{y}$ (r). The $\lambda_{3}$-term SOC gaps out all Dirac cones.
@@ C.4 Spin-orbit coupling
$H_{\rm SOC}=\lambda_{1}\sigma_{y}s_{z}(\cos\tfrac{k_{x}+k_{y}}{2}+\cos\tfrac{k_{x}-k_{y}}{2})+\lambda_{2}\sigma_{y}\tau_{z}s_{y}(\cos\tfrac{k_{x}+k_{y}}{2}+\cos\tfrac{k_{x}-k_{y}}{2})+\lambda_{3}\tau_{y}s_{x}\cos\tfrac{k_{z}}{2}$ (42)
@@
The SOC coupling terms have different effects on the electronic structure and the surface state, see Fig. 6 and Fig. 7. In summary, we conclude that SOC does not affect the emergence of altermagnetic surface states. Instead, it can even significantly increase the spin splitter effect because the spin degenerate bulk density is gapped out. SOC can alter the topology of the band structure, but the altermagnetic character of the surface remains intact.

### v2

@@ III.1 Non-magnetic part
$H$ $=2t\sigma_{x}(\cos\tfrac{k_{x}+k_{y}}{2}+\cos\tfrac{k_{x}-k_{y}}{2})+2t_{\perp}\tau_{x}\cos\tfrac{k_{z}}{2}$ (S22)
@@
$\bm{t}_{\tfrac{a}{2}\tfrac{c}{2}}$ $=I\{\tfrac{1}{2},0,\tfrac{1}{2}\}$ $H$ $=\sigma_{x}H\sigma_{x}$ (S23)
$\bm{t}_{\tfrac{b}{2}}$ $=I\{0,\tfrac{1}{2},0\}$ $H$ $=\tau_{x}H\tau_{x}$ (S24)
$\bm{t}_{\tfrac{a}{2}\tfrac{b}{2}\tfrac{c}{2}}$ $=I\{\tfrac{1}{2},\tfrac{1}{2},\tfrac{1}{2}\}$ $H$ $=\sigma_{x}\tau_{x}H\sigma_{x}\tau_{x}$ (S25)
which can give rise to accidental constraints for the band structure, as we show below. They can be broken by further neighbor hoppings, see Fig. S2(b). If we add
$\delta H=$ $2(t_{1}+t_{2})\sigma_{x}\tau_{x}\left[\cos\tfrac{k_{x}+k_{y}}{2}+\cos\tfrac{k_{x}-k_{y}}{2}\right]\cos\tfrac{k_{z}}{2}+2(t_{1}-t_{2})\sigma_{y}\tau_{y}\left[\sin\tfrac{k_{x}+k_{y}}{2}+\sin\tfrac{k_{x}-k_{y}}{2}\right]\sin\tfrac{k_{z}}{2}$ (S26)
@@
$\delta H$ $=2\sigma_{x}[(t_{3}+t_{4})\cos k_{z}(\cos\tfrac{k_{x}+k_{y}}{2}+\cos\tfrac{k_{x}-k_{y}}{2})]+2\sigma_{x}\tau_{z}[(t_{3}-t_{4})\sin k_{z}(\sin\tfrac{k_{x}+k_{y}}{2}-\sin\tfrac{k_{x}-k_{y}}{2})]$ (S27)
@@
$C_{2b}^{\prime}=C_{2b}\{\tfrac{1}{2},\tfrac{1}{2},\tfrac{1}{2}\}*I\{0,\tfrac{1}{2},0\}$ $=C_{2b}\{\tfrac{1}{2},0,\tfrac{1}{2}\}=\sigma_{x}$ (S28)
$C_{2b}^{\prime\prime}=C_{2b}\{\tfrac{1}{2},\tfrac{1}{2},\tfrac{1}{2}\}*I\{\tfrac{1}{2},\tfrac{1}{2},\tfrac{1}{2}\}$ $=C_{2b}\{0,0,0\}$ (S29)
@@ III.2 Magnetic part
$H_{M}=$ $2\Delta_{0}\sigma_{y}s_{z}(\sin\tfrac{k_{x}+k_{y}}{2}-\sin\tfrac{k_{x}-k_{y}}{2})+2\Delta_{1}\sigma_{z}\tau_{y}s_{z}\sin\tfrac{k_{z}}{2}$ (S30)
@@
$H_{M}^{\prime}=2\Delta_{2}\sigma_{y}\tau_{x}s_{z}(\sin\tfrac{k_{x}+k_{y}}{2}-\sin\tfrac{k_{x}-k_{y}}{2})\cos\tfrac{k_{z}}{2}]+2\Delta_{3}\sigma_{x}\tau_{y}s_{z}(\cos\tfrac{k_{x}+k_{y}}{2}-\cos\tfrac{k_{x}-k_{y}}{2})\sin\tfrac{k_{z}}{2}]$ (S32)
@@
III.3 Symmetry breaking perturbations
@@
$\delta H$ $=\sigma_{z}$ $(-,+,-)$ (S34)
$\delta H$ $=\tau_{z}$ $(-,-,+)$ (S35)
$\delta H$ $=\sigma_{z}\tau_{z}$ $(+,-,-)$ (S36)
@@
The perturbation $\tau_{z}$ therefore has the same symmetry as $u_{ab}$. However, it turns out that this perturbation is invariant under an accidental mirror $M_{b}^{\prime}=\sigma_{x}$ of the $k_{z}=0$ plane, and the nodal line is not gapped unless higher order, symmetry-allowed hoppings are included. Because of this, we consider another perturbation with the same symmetry $\sigma_{z}\tau_{z}\left[\cos\left(k_{x}+k_{y}\right)-\cos\left(k_{x}-k_{y}\right)\right]$, which is the hopping considered in the main text, and shown in Fig. S2 (d).
III.4 Spin-orbit coupling
Figure S3: Influence of SOC on the altermagnetic drumhead states. We consider each of the three SOC terms in Eq. S37, each time setting the other two to zero. (a-e) $\lambda_{1}=0.4$. Bulk 3D band structure (a), bulk spin-resolved spectral function (b,c), surface spin-resolved spectral function (d,e), and spin splitter angles $\alpha_{x},\alpha_{y}$ (f). In the presence of $\lambda_{1}$-term SOC the spin-space group notation breaks down such that $C_{2}||M_{a}]$ is not a symmetry anymore. (g-k) $\lambda_{2}=0.2$. Bulk 3D band structure (g), bulk spin-resolved spectral function (h,i), surface spin-resolved spectral function (j,k), and spin splitter angles $\alpha_{x},\alpha_{y}$ (l). The $\lambda_{2}$-term SOC gaps out the Dirac nodal line everywhere except at the BZ boundary, where 2 Dirac cones persist and spin remains a good quantum number. The topologically unprotected, altermagnetic surface state persists. (m-q) $\lambda_{3}=0.2$. Bulk 3D band structure (m), bulk spin-resolved spectral function (n,o), surface spin-resolved spectral function (p,q), and spin splitter angles $\alpha_{x},\alpha_{y}$ (r). The $\lambda_{3}$-term SOC gaps out the Dirac nodal line everywhere. The topologically unprotected, altermagnetic surface state persists.
Figure S4: Influence of SOC on the altermagnetic Fermi-arc state. We consider each of the three SOC terms in Eq. S37, each time setting the other two to zero. (a-e) $\lambda_{1}=0.4$. Bulk 3D band structure (a), bulk spin-resolved spectral function (b,c), surface spin-resolved spectral function (d,e), and spin splitter angles $\alpha_{x},\alpha_{y}$ (f). In the presence of $\lambda_{1}$-term SOC the spin-space group notation breaks down such that $C_{2}||M_{a}]$ is not a symmetry anymore. (g-k) $\lambda_{2}=0.2$. Bulk 3D band structure (g), bulk spin-resolved spectral function (h,i), surface spin-resolved spectral function (j,k), and spin splitter angles $\alpha_{x},\alpha_{y}$ (l). The $\lambda_{2}$-term SOC gaps out the Dirac cones along the $y$-axis.The topologically unprotected, altermagnetic surface state persists. (m-q) $\lambda_{3}=0.2$. Bulk 3D band structure (m), bulk spin-resolved spectral function (n,o), surface spin-resolved spectral function (p,q), and spin splitter angles $\alpha_{x},\alpha_{y}$ (r). The $\lambda_{3}$-term SOC gaps out all Dirac cones.
@@ III.4 Spin-orbit coupling
$H_{\rm SOC}=\lambda_{1}\sigma_{y}s_{z}(\cos\tfrac{k_{x}+k_{y}}{2}+\cos\tfrac{k_{x}-k_{y}}{2})+\lambda_{2}\sigma_{y}\tau_{z}s_{y}(\cos\tfrac{k_{x}+k_{y}}{2}+\cos\tfrac{k_{x}-k_{y}}{2})+\lambda_{3}\tau_{y}s_{x}\cos\tfrac{k_{z}}{2}$ (S37)
@@
The SOC coupling terms have different effects on the electronic structure and the surface state, see Fig. S3 and Fig. S4. In summary, we conclude that SOC does not affect the emergence of altermagnetic surface states. Instead, it can even significantly increase the spin splitter effect because the spin degenerate bulk density is gapped out. SOC can alter the topology of the band structure, but the altermagnetic character of the surface remains intact.